A baseball team played 143 complete games last season. They had 29 more wins than losses. How many games did the team lose?
Answer by algebrahouse.com(1659)

"A baseball team played 143 complete games last season. They had 29 more wins than losses. How many games did the team lose?"

x = number of losses
x + 29 = number of wins {they had 29 more wins than losses}

x + x + 29 = 143 {played 143 total games}
2x + 29 = 143 {combined like terms}
2x = 114 {subtracted 29 from both sides}
x = 57 {divided both sides by 2}

the team had 57 losses

For more help from me, visit: www.algebrahouse.com